Suppose a consumers preferences can be represented by the utility function U(x,y)=X2Y .in the table below, identify 4 point s on the consumer s indifference curve where U=10000, find x,y


1
Expert's answer
2021-09-06T07:06:25-0400

Considering the following utility function’ 

Utility (U) = X2

 

Now when the utility is 10000, 

10000 = X2

"Or, X= \sqrt{\frac{10000}{X}}"

 

The required equation is "X= \sqrt{\frac{10000}{X}}"